A package of 6 cellphones contains 4 that are slightly defective. Elise bought 4 of these CP's at random. Let the random variable be the number of non-defective CP's.





a) List the elements of the sample space using D and N for defective and non defective cell phones respectively.

1
Expert's answer
2022-05-22T23:04:57-0400

a)


"S=\{DDDD, DDDN, DDND, DNDD,"

"NDDD,DDNN,DNDN, DNND,"


"NDND, NDDN, NNDD\}"



b)


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & 0 & 1 & 2 \\ \hline
 p(x) & 1/11 & 4/11 & 6/11 \\
 
\end{array}"

c)


"\mu=E(X)=\dfrac{1}{11}(0)+\dfrac{4}{11}(1)+\dfrac{6}{11}(2)=\dfrac{16}{11}"

d)


"E(X^2)=\dfrac{1}{11}(0)^2+\dfrac{4}{11}(1)^2+\dfrac{6}{11}(2)^2=\dfrac{28}{11}"

"Var(X)=\sigma^2=E(X^2)-(E(X))^2=\dfrac{28}{11}-(\dfrac{16}{11})^2"

"=\dfrac{52}{121}"

e)


"\sigma=\sqrt{\sigma^2}=\sqrt{\dfrac{52}{121}}=\dfrac{2\sqrt{13}}{11}"